How could you prepare 250 ml 0.20M using only a solution of 1.0M NaCl and water?


1
Expert's answer
2021-01-28T05:48:52-0500

According to the equation:

C1 × V1 = C2 × V2

where C1 - the initial concentration, V1 - the volume of the initial solution, C2 - the final volume, V2 - the volume of the final solution.

From here, the volume of the initial solution required:

V1 = C2 × V2 / C1 = 250 ml × 0.20 M / 1.0 M = 50 ml

As a result, to prepare 250 ml of 0.20M NaCl, take 50 ml of 1.0M NaCl solution and add 200 ml of water.